OpenStudy (anonymous):

Factor completely 40xy2 - 36x2y. 4xy(9y - 10x) 4xy(10y - 9x) Prime 4xy(10x - 9y) @Whitemonsterbunny17 @zaibali.qasmi

OpenStudy (anonymous):

@AravindG

OpenStudy (anonymous):

@amistre64

whitemonsterbunny17 (whitemonsterbunny17):

Factoring... Sorry, but I suck at that. :s @AravindG might be able to help tho c:

OpenStudy (anonymous):

thx for trying

OpenStudy (aravindg):

Take out the common terms.

OpenStudy (anonymous):

how does he do it @gregmb1999

whitemonsterbunny17 (whitemonsterbunny17):

IDK!! He's a ghost!!! O.O

OpenStudy (anonymous):

he sneaks and says the answers and then blocks me

OpenStudy (anonymous):

so is it A

OpenStudy (anonymous):

@gregmb1999

OpenStudy (anonymous):

is it B or A @AravindG

OpenStudy (anonymous):

HAI ^.^ @gregmb1999 i have a question are you a ghosy for realz >.<

OpenStudy (aravindg):

Open up the brackets using distributive property and realize the wonder for yourself.

OpenStudy (anonymous):

okay so 4xy(10y-9x)?

whitemonsterbunny17 (whitemonsterbunny17):

@gregmb1999, please do not give out answers, as it is against the CoC (Code of Conduct). You may help guide the user to the answer as @AravindG is doing, but please do not give direct answers.

OpenStudy (anonymous):

@gregmb1999 HAI could you help me or @AravindG im kind of understanding

OpenStudy (anonymous):

yes I agree @Whitemonsterbunny17

OpenStudy (aravindg):

Consider a(b+c)=ab+ac 4xy(9y - 10x)=? 4xy(10y - 9x)=? Can't you do it now?

OpenStudy (anonymous):

\(4xy(9y - 10x)=36xy^2-40x^2y\) and \(4xy(10y - 9x)=40xy^2-36x^2y\) is it the second one?

OpenStudy (anonymous):

@amistre64 and @AravindG and @Whitemonsterbunny17

OpenStudy (aravindg):

Yes :)

OpenStudy (aravindg):

Good work!

OpenStudy (anonymous):

thx ;) \(\huge\ddot\smile\ ~\checkmark\)

OpenStudy (aravindg):

yw :)